how do i do -2/5(10x-25)=8x+4??
Answer by xcentaur(357)

-2/5(10x-25)=8x+4
Multiplying throughout by 5 we get,
-2(10x-25)=5(8x+4)
Opening the brackets,
-20x+50=40x+20
Taking all the x to the left sides and all the numbers to the right,
-20x-40x=20-50
-60x=-30
60x=30
x=30/60
x=1/2

Thats your answer
Hope this helps,
Prabhat